OpenStudy (anonymous):

What are the foci of the hyperbola given by the equation 16x2 - 25y2 - 64x - 50y - 361 = 0?

OpenStudy (anonymous):

The hyperbola has the simplified equation (4x-8)²/20² - (5y+5)²/ 20² =1 4²(x-2)²/20² - 5²(y+1)²/ 20² =1 (x-2)²/5² -(y+1)²/4² =1 in new coordinates X=x-2 Y=y+1 we have X²/5² - Y²/4² =1 The foci are (-c,0), (c,0) where c²=a²+b² a²=25 b²=16 c²=41 c=√41 F1=(-√41 ,0) F2=(√41,0) In original coordinates f1= (2-√41 , -1) f2=(2+√41 , -1)

OpenStudy (anonymous):

wow that was quick and thorough!

OpenStudy (anonymous):

YOU"RE AWSOME